solve each system by graphing.
x - 2y = 8
3x - 2y = 12

Answer by ryanman(19)

-(x - 2y = 8)
3x - 2y = 12
----------
2x =4
x=2
now substitute
3(2) -2y =12
6-2y=12
-6=2y
y=-3